OpenStudy (anonymous):

To help find the velocity of particles requires the evaluation of the indenite integral of the acceleration function, a(t), i.e. v = Z ∫a(t)dt Evaluate the following indenite integrals: (a) ∫??lntt4dt (b) ∫??(t2+1)sin9tdt

OpenStudy (anonymous):

To be absolutely clear, is that \[\int t^4 \ln t~dt~~\text{and}~~\int (t^2+1)\sin9t~dt~~?\]

OpenStudy (anonymous):

a) \[\int\limits_{?}^{?} \frac{ lnt }{ t^4 } dt\]

OpenStudy (anonymous):

b) yes

OpenStudy (anonymous):

I think integrating by parts will work for both: \[\begin{matrix}u=\ln t&&&dv=\frac{dt}{t^4}\\ du=\frac{dt}{t}&&&v=-\frac{1}{3t^3}\end{matrix}\] Then \[\int \frac{\ln t}{t^4}~dt=-\frac{\ln t}{3t^3}+\int\frac{dt}{t^4}\] Oof, the second one is significantly more tedious...

OpenStudy (anonymous):

could you show me the full working out at all please?

OpenStudy (anonymous):

That would be doing your work for you, so no, but I'm willing to start you off. I think I've given you enough to tackle the first. Speaking of which, there's a slight mistake, it should be \[\int \frac{\ln t}{t^4}~dt=-\frac{\ln t}{3t^3}+\color{red}{\frac{1}{3}}\int\frac{dt}{t^4}\]

OpenStudy (anonymous):

For the second one, you integrate by parts twice: \[\begin{matrix}f_1=t^2+1&&&dg_1=\sin9t\\ df_1=2t~dt&&&g_1=-\frac{1}{9}\cos9t \end{matrix}\] So \[\begin{align*}\int(t^2+1)\sin9t~dt&=-\frac{1}{9}(t^2+1)\cos9t+\frac{2}{9}\int t\cos9t~dt \end{align*}\] By parts again, letting \[\begin{matrix}f_2=t&&&dg_2=\cos9t\\ df_2=dt&&&g_2=\frac{1}{9}\sin9t \end{matrix}\] Carry on...

OpenStudy (anonymous):

Perfect :) @SithsAndGiggles

OpenStudy (anonymous):

Thanks :3

OpenStudy (anonymous):

what was your final answer? sorry thank you for your help, just checking I have done it right.

OpenStudy (anonymous):

@SithsAndGiggles